find the measure of an angle if it's measure is 9 degrees less than 8 times its complement
Answer by rapaljer(4551)

Let x= the complement of the angle
8x - 9 = the angle itself

x + 8x -9 = 90
9x = 99
x=11 the complement of the angle
8x-9= 8*11 -9=88-9 =79 degrees = the angle

Check: 79+11=90

R^2